# How Far Is Video Generation from World Model: A Physical Law Perspective

### v1

@@ 4 Combinatorial Generalization
In Section 3, video generation models failed to reason in OOD scenarios. This is understandable—deriving precise physical laws from data is difficult for both humans and models. For example, it took scientists centuries to formulate Newton’s three laws of motion. However, even a child can intuitively predict outcomes in everyday situations by combining elements from past experiences. This ability to combine known information to predict new scenarios is called combinatorial generalization. In this section, we evaluate the combinatorial abilities of diffusion-based video models.
@@ 4.1 Combinatorial Physical Scenarios
We selected the PHYRE simulator (Bakhtin et al., 2019) as our testbed—a 2D environment involves multiple objects to free fall then collide with each other, forming complex physical interactions. It features diverse object types, including balls, jars, bars, and walls, which can be either fixed or dynamic. This enables complex interactions such as collisions, parabolic trajectories, rotations, and friction to occur simultaneously within a video. Despite this complexity, the underlying physical laws are deterministic, allowing the model to learn the laws and predict unseen scenarios.
Training Data. There are eight types of objects considered, including two dynamic gray balls, a group of fixed black balls, a fixed black bar, a dynamic bar, a group of dynamic standing bars, a dynamic jar, and a dynamic standing stick. Each task contains one red ball and four randomly chonsen objects from the eight types, resulting in unique templates. See Figure 4 for examples.
Each template was initialized with random sizes and positions for four objects, generating 100K videos to cover a range of possible scenarios. To explore the model’s combinatorial ability and scaling effects, we structured the training data at three levels: a minimal set of 6 templates (0.6M videos) that includes all types of two-object interactions among the eight object types, and larger sets with 30 and 60 templates (3M/6M videos), with the 60-template set nearly covering the entire template space. The minimal training set places the highest demand on the model’s ability for compositional generalization.
@@
Models. The first frame is used as the conditioning for video generation since the inital objects are static. We found that smaller models like DiT-S struggled with complex videos, so we primarily used DiT-B and DiT-XL. All models were trained for long 1000K gradient steps on 64 Nvidia A100 GPUs with a batch size of 256, ensuring near convergence. To better capture the complexity of physical events, we increased the resolution to 256x256 with 32 frames.
Evaluation Metrics. We use several metrics to assess the fidelity of generated videos compared to the ground truth. Frechet Video Distance (FVD) (Unterthiner et al., 2018) calculates feature distances between generated and real videos using features from Inflated-3D ConvNets (I3D) pretrained on Kinetics-400 (Carreira & Zisserman, 2017). SSIM and PSNR (Wang et al., 2004) are pixel-level metrics: SSIM evaluates brightness, contrast, and structural similarity, while PSNR measures the ratio between peak signal and mean squared error, both averaged across frames. LPIPS (Zhang et al., 2018) gauges perceptual similarity between image patches. We include human evaluations, reporting the abnormal ratio of generated videos that violate physical laws assessed by humans.
4.2 Main Results
{in-template result} / {out-of-template result}.
@@ 4.2 Main Results
It requires higher resolution, much more training iterations, and larger model sizes to perform well on this task due to increased complexity. Consequently, we are unable to conduct a comprehensive sweep of all data and model size combinations as in Section 3. Therefore, we start with the largest model, DiT-XL, to study data scaling behavior for combinatorial generalization. As shown in Table 2, when the number of templates increases from 6 to 60, all metrics improve on the out-of-template testing sets. Notably, the abnormal rate for human evaluation significantly reduces from 67% to 10%. Conversely, the model trained with 6 templates achieves the best SSIM, PSNR, and LPIPS scores on the in-template testing set. This can be explained by the fact that each training example in the 6-template set is exposed ten times more frequently than those in the 60-template set, allowing it to better fit the in-template tasks associated with template 6. Furthermore, we conducted an additional experiment using a DiT-B model on the full 60 templates to verify the importance of model scaling. As expected, the abnormal rate increases to 24%. These results suggest that both model capacity and coverage of the combination space are crucial for combinatorial generalization. This insight implies that scaling laws for video generation should focus on increasing combination diversity, rather than merely scaling up data volume. The video generation visualizations of our models can be found in Figure 17 and Figure 18.
@@ 5.1 Understanding Generalization from Interpolation and Extrapolation
The generalization ability of a model roots from its interpolation and extrapolation capability (Xu et al., 2020; Balestriero et al., 2021). In this section, we design experiments to explore the limits of these abilities for a video generation model. We design datasets which delibrately leave out some latent values, i.e. velocity. After training, we test model’s prediction on both seen and unseen scenarios. We mainly focus on uniform motion and collision processes.
Uniform Motion. We create a series of training sets, where a certain range of velocity is absent. Each set contains 200K videos to ensure fairness. As shown in Figure 5 (1)-(2), with a large gap in the training set, the model tends to generate videos where the velocity is either high or low to resemble training data when initial frames show middle-range velocities. We find video generation model’s OOD accuracy is closely related to the size of gap, as seen in Figure 5 (3), when the gap is reduced, the model correctly interpolates for most of OOD data. Moreover, as shown in Figure 5 (4) and (5), when a subset of the missing range is reintroduced (without increasing data amount), the model exhibits stronger interpolation abilities.
Collision involves multiple variables, which is more challenging since the model has to learn a two-dimensional non-linear function. Specifically, we exclude one or more square regions from the training set of initial velocities for two balls and then assess the velocity prediction error after the collision. For each velocity point, we sample a grid of radius parameters to generate multiple video cases and compute the average error. As shown in Figure 6 (1)-(2), an interesting phenomenon happens. The video generation model’s extrapolation error demonstate an intringuing discrepancy among OOD points: For the OOD velocity combinations that lie within the convex hull of the training set, i.e., the internal red squares in the yellow region, the model generalizes well. However, the model experiences large errors when the latent values lies in exterior space of training set’s convex hull.
@@ 5.2 Memorization or Generalization
We train our model on uniform motion videos with velocities , using the first three frames as input conditions. Two training sets are used: Set-1 only contains balls moving from left to right, while Set-2 includes movement in both direction, by using horizontal flipping at training time. At evaluation, we focus on low-speed balls (), which were not present in the training data. As shown in Figure 7, the Set-1 model generates videos with only positive velocities, biased toward the high-speed range. In contrast, the Set-2 model occasionally produces videos with negative velocities, as highlighted by the green circle. For instance, a low-speed ball moving from left to right may suddenly reverse direction after its condition frames. This could occur since the model identifies reversed training videos as the closest match for low-speed balls. This distinction between the two models suggests that the video generation model is influenced by “deceptive” examples in the training data. Rather than abstracting universal rules, the model appears to rely on memorization, and case-based imitation for OOD generalization.
@@ 5.3 How Does Diffusion Model Retrieve Data?
The other three pairs are presented in Figure 9. For velocity vs. size, the combinatorial generalization performance is surprisingly good. The model effectively maintains the initial size and velocity for most test cases beyond the training distribution. However, a slight preference for size over velocity is noted, particularly with extreme radius and velocity values (top left and bottom right in Figure 9 (1)). In Figure 9 (2), color can be combined with size most of the time. Conversely, for color vs. velocity in Figure 9 (3), high-speed blue balls and low-speed red balls are used for training. At test time, low-speed blue balls appear much faster than their conditioned velocity. No ball in the testing set changes its color, indicating that color is prioritized over velocity. Based on the above analysis, we conclude that the prioritization order is as follows: color > size > velocity > shape.
5.4 How Does Complex Combinatorial Generalization Happen?
@@
Spatial composition. As given by Figure 11 (left side) in the appendix, the training data contains two distinct types of physical events. One type involves a blue square moving horizontally with a constant velocity while a red ball remains stationary. In contrast, the other type depicts a red ball moving toward and then bouncing off a wall while the blue square remains stationary. At test time, when the red ball and the blue square are moving simultaneously, the learned model is able to generate the scenario where the red ball bounces off the wall while the blue square continues its uniform motion.
Temporal combination. As illustrated on the right side of Figure 11, when the training data includes distinct physical events—half featuring two balls colliding without bouncing and the other half showing a red ball bouncing off a wall—the model learns to combine these events temporally. Consequently, during evaluation, when the balls collide near the wall, the model accurately predicts the collision and then determines that the blue ball will rebound off the wall with unchanged velocity.
With these attribute, spatial, and temporal combinatorial patterns, the video generation model can identify basic physical events in the training set and combine them across attributes, time, and space to generate videos featuring complex chains of physical events.
@@ 5.5 Is Video Sufficient for Complete Physics Modeling?
For a video generation model to function as a world model, the visual representation must provide sufficient information for complete physics modeling. In our experiments, we found that visual ambiguity leads to significant inaccuracies in fine-grained physics modeling. For example, in Figure 10, it is difficult to determine if a ball can pass through a gap based on vision alone when the size difference is at the pixel level, leading to visually plausible but incorrect results. Similarly, visual ambiguity in a ball’s horizontal position relative to a block can result in different outcomes. These findings suggest that relying solely on visual representations, may be inadequate for accurate physics modeling.
@@ 6 Related Works
Video generation. Open-set video generation is mostly based on diffusion models (Rombach et al., 2022; Ho et al., 2022b; a; He et al., 2024) or auto-regressive models (Yu et al., 2023a; b; Kondratyuk et al., 2023). These models often require a pretrained image or video VAE (Kingma, 2013; Van Den Oord et al., 2017) for data compression to improve computational efficiency. Some approaches leverage pretrained Text-to-Image (T2I) models for zero-shot (Khachatryan et al., 2023; Zhang et al., 2023b) or few-shot (Wu et al., 2023) video generation. Additionally, Image-to-Video(I2V) generation (Zeng et al., 2024; Girdhar et al., 2023; Zhang et al., 2023a) shows that video quality improves substantially when conditioned on an image. The Diffusion Transformer (DiT) (Peebles & Xie, 2023) demonstrates better scaling behavior than U-Net (Ronneberger et al., 2015) for T2I generation. Sora (Brooks et al., 2024) leverages the DiT architecture to directly operate on spacetime patches of video and image latent codes. Our model follows Sora’s architecture and conceptually aligns with I2V generation, relying on image(s) for conditioning instead of text prompts.
@@
World model. World models (Ha & Schmidhuber, 2018) aim to learn models that can accurately predict how an environment evolves after some actions are taken. Previously, they often operated in an abstracted space and were used in reinforcement learning (Sutton, 2018) to enable planning (Silver et al., 2016; Schrittwieser et al., 2020) or facilitate policy learning through virtual interactions (Hafner et al., 2019; 2020). With the advancement of generative models, world models can now directly work with visual observations by employing a general framework of conditioned video generation. For example, in autonomous driving (Hu et al., 2023; Jia et al., 2023; Gao et al., 2024; Zheng et al., 2024), the condition is the driver’s operations, while in robot world models (Yang et al., 2023; Black et al., 2023; 1xw, 2024), the condition is often the control signals. Genie (Bruce et al., 2024) instead recovers the conditions from video games in an unsupervised learning manner. In our physical law discovery setting, it does not require a per-step action/condition since the physical event is determined by the underlying laws once an initial state is specified. See more related works in Section A.1.
7 Conclusion
@@
Video generation is believed as a promising way towards scalable world models. However, its capability to learn physical laws from visual observations has not yet been verified. We conducted the first systematic study in this area by examining its generalization performance across three typical scenarios: in-distribution, out-of-distribution (OOD), and combinatorial generalization. The findings indicate that scaling alone cannot address the OOD problem, although it does enhance performance in other scenarios. Our in-depth analysis suggests that video model generalization relies more on referencing similar training examples rather than learning universal rules. We observed a prioritization order of color > size > velocity > shape in this "case-based" behavior. In conclusion, our study suggests that naively scaling is insufficient for video generation models to discover fundamental physical laws.
Appendix A Appendix
A.1 More related works

### v2

@@ 4 Combinatorial Generalization
In Section 3, video generation models failed to reason in the OOD scenarios. This is understandable–deriving precise physical laws from data is difficult for both humans and models. For example, it took scientists centuries to formulate Newton’s three laws of motion. However, even a child can intuitively predict outcomes in everyday situations by combining elements from past experiences. This ability to combine known information to predict new scenarios is called combinatorial generalization. Now, we evaluate the combinatorial abilities of diffusion-based video models.
@@ 4.1 Combinatorial Physical Scenarios
We selected the PHYRE simulator (Bakhtin et al., 2019) as our testbed—a 2D environment involves multiple objects that fall and then collide with each other, forming complex physical interactions. It features various types of objects, including balls, jars, bars, and walls, which can be either fixed or dynamic. This enables complex interactions such as collisions, parabolic trajectories, rotations, and friction to occur simultaneously within a video. Despite this complexity, the underlying physical laws are deterministic, allowing the model to learn the laws and predict unseen scenarios.
Training Data. There are eight types of objects are considered, including two dynamic gray balls, a group of fixed black balls, a fixed black bar, a dynamic bar, a group of dynamic standing bars, a dynamic jar and a dynamic standing stick. Each task contains one red ball and four randomly chonsen objects from the eight types, resulting in unique templates. See Figure 4 for examples. Each object has a distinct visual appearance and color, allowing the model to infer object type and physical attributes by learning associations from training data. Since all objects start at rest with zero initial velocity, we provide the first frame so that the model can infer their attributes and use this information to predict future events.
Each template was initialized with random sizes and positions for four objects, generating 100K videos to cover a range of possible scenarios. To explore the combinatorial ability and scaling effects of the model, we structured the training data at three levels: a minimal set of 6 templates (0.6M videos) that includes all types of interaction of two objects between the eight types of objects, and larger sets with 30 and 60 templates (3M/6M videos), with the set of 60 templates almost covering the entire template space. The minimal training set places the greatest demand on the model’s ability for compositional generalization.
@@
Models. The first frame is used as the conditioning for video generation since the initial objects are static. We found that smaller models such as DiT-S struggled with complex videos, so we mainly used DiT-B and DiT-XL. All models were trained for 1000K gradient steps on 64 Nvidia A100 GPUs with a batch size of 256, ensuring convergence. To better capture the complexity of physical events, we increased the resolution to 256x256 with 32 frames.
Evaluation Metrics. In the ID/OOD setting, scenes are simple (e.g., 1–2 colored balls on a plain background), enabling reliable position and velocity estimation via pixel averaging and frame differencing. In contrast, the combinatorial setting includes many irregularly shaped and similarly colored objects, making pixel assignments ambiguous and position estimation unreliable. We also tested methods such as Hough circle detection and pretrained object detectors, but they resulted in significant errors. Given these challenges, we excluded velocity as a metric.
Instead, we rely on a combination of objective video fidelity metrics and human evaluation. We use several metrics to assess the fidelity of the generated videos compared to ground truth. Frechet Video Distance (FVD) (Unterthiner et al., 2018) calculates the feature distances between generated and real videos using features from Inflated-3D ConvNets (I3D) pretrained on Kinetics-400 (Carreira & Zisserman, 2017). SSIM and PSNR (Wang et al., 2004) are pixel-level metrics: SSIM evaluates brightness, contrast, and structural similarity, while PSNR measures the ratio between peak signal and mean squared error, both averaged across frames. LPIPS (Zhang et al., 2018) gauges the perceptual similarity between image patches. Since pixel-level metrics may not fully capture physical plausibility, we also include human evaluations, reporting the abnormal ratio of videos generated that violate physical laws assessed by humans. This combined approach offers a comprehensive evaluation for physical correctness in the complex setting.
4.2 Scaling Law Observed for Combinatorial Generalization
@@
It requires higher resolution, much more training iterations, and larger model sizes to perform well on this task due to increased complexity. Consequently, we are unable to conduct a comprehensive sweep of all data and model size combinations as in Section 3. Therefore, we start with the largest model, DiT-XL, to study data scaling behavior for combinatorial generalization. As shown in Table 2, when the number of templates increases from 6 to 60, all metrics improve on the out-of-template testing sets. Notably, the abnormal rate for human evaluation significantly reduces from 67% to 10%. Conversely, the model trained with 6 templates achieves the best SSIM, PSNR, and LPIPS scores on the in-template testing set. This can be explained by the fact that each training example in the 6-template set is exposed ten times more frequently than those in the 60-template set, allowing it to better fit the in-template tasks associated with template 6. Furthermore, we conducted an additional experiment using a DiT-B model on the full 60 templates to verify the importance of model scaling. As expected, the abnormal rate increases to 24%. These results suggest that both model capacity and coverage of the combination space are crucial for combinatorial generalization. This insight implies that scaling laws for video generation should focus on increasing combination diversity, rather than merely scaling up data volume. Some generated videos can be found in Figure 21, Figure 22, Figure 23, and Figure 24.
@@ 5.1 Understanding Generalization from Interpolation and Extrapolation
The generalization ability of a model originates from its interpolation and extrapolation ability (Xu et al., 2020; Balestriero et al., 2021). In this section, we design experiments to explore the limits of these abilities for a video generation model. We design datasets which delibrately leave out some latent values, i.e. velocity. After training, we test model’s prediction on both seen and unseen scenarios. We mainly focus on uniform motion and collision processes.
Uniform Motion. We create a series of training sets, where a certain range of velocity is absent. Each set contains 200K videos to ensure fairness. As shown in Figure 5 (1)-(2), with a large gap in the training set, the model tends to generate videos where the velocity is high or low to resemble the training data when initial frames show middle-range velocities. We find that the video generation model’s OOD accuracy is closely related to the size of the gap, as seen in Figure 5 (3), when the gap is reduced, the model correctly interpolates for most of the OOD data. Moreover, as shown in Figure 5 (4) and (5), when a subset of the missing range is reintroduced (without increasing the amount of data), the model exhibits stronger interpolation abilities. In total, In Section 3, the evaluation is strictly OOD extrapolation. In this setting, simply increasing the amount of data or model size does not significantly improve performance due to the difficulty of truly grasping the physical law and true extrapolation. In contrast, Figure 5 reveals a transition from extrapolation to interpolation. As the gap between the adjacent training regions narrows, model performance on the test set improves, reflecting the model’s stronger ability to interpolate.
Collision. It involves multiple variables and is more challenging since the model has to learn a two-dimensional nonlinear function. Specifically, we exclude one or more square regions from the training set of initial velocities for two balls and then assess the velocity prediction error after the collision. For each velocity point, we sample a grid of radius parameters to generate multiple video cases and compute the average error. As shown in Figure 6 (1)-(2), the video generation model’s extrapolation error induces an intringuing discrepancy among OOD points: For the OOD velocity combinations that lie within the convex hull of the training set, i.e., the internal red squares in the yellow region, the model generalizes well. However, the model experiences large errors when the latent values lie in the exterior space of the convex hull of the training set.
@@ 5.2 Memorization or Generalization
We train our model on uniform motion videos with velocities , using the first three frames as input conditions. Two training sets are used: Set-1 only contains balls moving from left to right, while Set-2 includes movement in both directions, by using horizontal flipping at training time. In evaluation, we focus on low-speed balls (), which were not present in the training data. As shown in Figure 7, the Set-1 model generates videos with only positive velocities, biased toward the high-speed range. In contrast, the Set-2 model occasionally produces videos with negative velocities, as highlighted by the green circle. For instance, a low-speed ball moving from left to right may suddenly reverse direction after its condition frames. This could occur since the model identifies reversed training videos as the closest match for low-speed balls. This distinction between the two models suggests that the video generation model is influenced by “deceptive” examples in the training data. Rather than abstracting universal rules, the model appears to rely on memorization and case-based imitation for OOD generalization.
@@ 5.3 How Does Diffusion Model Retrieve Data?
The other three attribute pairs are presented in Figure 9. For the continuous attributes of velocity and size, we selected two distinct values at the extreme boundaries to create binary comparisons: 1.0 and 4.0 for velocity, and 0.7 and 1.4 for size. For the comparison of color vs. size, the training set consists of: (1) red, small balls and (2) blue, large balls. During evaluation, across 202 evaluation cases, the size was consistently transformed while the color remained fixed, clearly demonstrating that the color size. Similar observations are evident in Figure 9 (2) and (3), leading to the conclusion that color velocity and size velocity. Based on the above analysis, we establish the prioritization order as color size velocity shape.
5.4 Complex Combinatorial Generalization
@@
Spatial composition. As given by Figure 11 (left side) in the Appendix, the training data contain two distinct types of physical events. One type involves a blue square moving horizontally with a constant velocity while a red ball remains stationary. In contrast, the other type depicts a red ball moving toward and then bouncing off a wall, while the blue square remains stationary. At test time, when the red ball and the blue square are moving simultaneously, the model is able to correctly generate the movement of both balls.
Temporal combination. As illustrated on the right side of Figure 11, when the training data includes distinct physical events—half featuring two balls colliding without bouncing and the other half showing a red ball bouncing off a wall—the model learns to combine these events temporally. Consequently, when the balls collide near the wall, the model accurately predicts the collision and the blue ball will rebound off the wall with unchanged velocity.
Therefore, the video generation model can identify basic physical events in the training set and combine them across attributes, time, and space to generate videos featuring complex chains of physical events.
@@ 5.5 Is Video Sufficient for Complete Physics Modeling?
For a video generation model to function as a world model, the visual representation must provide sufficient information for complete physics modeling. In our experiments, we found that visual ambiguity leads to significant inaccuracies in fine-grained physics modeling. For example, in Figure 10, it is difficult to determine if a ball can pass through a gap based on vision alone when the size difference is at the pixel level, leading to visually plausible but incorrect results. Similarly, visual ambiguity in the horizontal position of a ball relative to a block can result in different outcomes. These findings suggest that relying solely on visual representations, may be inadequate for accurate physics modeling.
@@ 6 Related Works
Video generation. Open-set video generation is mainly based on diffusion models (Rombach et al., 2022; Ho et al., 2022b, a; He et al., 2024) or auto-regressive models (Yu et al., 2023a, b; Kondratyuk et al., 2023). These models often require a pre-trained image or video VAE (Kingma, 2013; Van Den Oord et al., 2017) for data compression to improve computational efficiency. Some approaches leverage pre-trained text-to-image (T2I) models for zero-shot (Khachatryan et al., 2023; Zhang et al., 2023b) or few-shot (Wu et al., 2023) video generation. Furthermore, Image-to-Video (I2V) generation (Zeng et al., 2024; Girdhar et al., 2023; Zhang et al., 2023a) shows that video quality improves substantially when conditioned on an image. The Diffusion Transformer (DiT) (Peebles & Xie, 2023) demonstrates a better scaling behavior than U-Net (Ronneberger et al., 2015) for T2I generation. Sora (Brooks et al., 2024) leverages the DiT architecture to directly operate on space-time patches of latent video and image codes. Our model follows Sora’s architecture and conceptually aligns with I2V generation, relying on image(s) for conditioning instead of text prompts. There are several previous work evaluating the physical understanding of large video models (Bansal et al., 2024; Sun et al., 2024; Liao et al., 2024) and benchmarks evaluating large video models more broadly (Huang et al., 2024; Liu et al., 2024).
@@
World model. World models (Ha & Schmidhuber, 2018) aim to learn models that can accurately predict how an environment evolves after some actions are taken. Previously, they often operated in an abstracted space and were used in reinforcement learning (Sutton, 2018) to enable planning (Silver et al., 2016; Schrittwieser et al., 2020) or facilitate policy learning through virtual interactions (Hafner et al., 2019, 2020; Yue et al., 2023). With the advancement of generative models, world models can now directly work with visual observations by employing a general framework of conditioned video generation. For example, in autonomous driving (Hu et al., 2023; Jia et al., 2023; Gao et al., 2024; Zheng et al., 2024), the condition is the driver’s operations, while in robot world models (Yang et al., 2023; Black et al., 2023; 1xw, 2024), the condition is often the control signals. Genie (Bruce et al., 2024) instead recovers the conditions from video games in an unsupervised way. In our physical law discovery setting, it does not require a per-step action/condition since the physical event is determined by the underlying laws once an initial state is specified. See more related works in Appendix A.
7 Conclusion and Discussion
@@
Conclusion. Video generation is believed to be a promising way toward scalable world models. However, its capability to learn physical laws from visual observations has not yet been verified. We conducted the first systematic study in this area by examining its generalization performance in three typical scenarios: in-distribution, out-of-distribution (OOD), and combinatorial generalization. The findings indicate that scaling alone cannot address the OOD problem, although it does enhance performance in other scenarios. Our in-depth analysis suggests that the generalization of the video model is based more on referencing similar training examples than learning universal rules. We observed a prioritization order of the color size velocity shape in this “case-based” behavior. In conclusion, our study suggests that naive scaling is insufficient to discover physical laws.
Disccusion. The insights from our study have practical implications for real-world scenarios. Real-world video data is inherently vast, continuous, and high-dimensional, making it much harder to comprehensively cover compared to pure language data. For instance, in real-world robotics, factors such as object velocity, joint configurations, camera perspectives, background noise, and task goals vary across a continuous and combinatorially large space. Our results suggest a path forward for addressing these challenges - increasing the combinatorial diversity of training data leads to significantly better performance in physical video modeling. Moreover, our findings imply that model scaling is most effective when supported by a sufficiently diverse training distribution. This has direct implications for fututre work to design both datasets and models that can generalize robustly in complex, real-world environments.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix Contents
[appendix] 1
Appendix A More related works